# Machine Learning to Predict Spectral Anisotropy in Valence-to-Core X-ray Emission Spectroscopy

### v1

@@ 5 Acknowledgments
JJK and JJR acknowledge support from the Theory Center for Materials and Energy Spectroscopies (TIMES) at SLAC funded by DOE BSE Contract DE-AC02-76SF00515. CAC was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2140004. GTS was supported by funding from the U.S. Department of Energy in the Nuclear Energy University Program under Contract No. DE-NE0009158. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or the U.S. Department of Energy. This work was partly facilitated through the use of advanced computational infrastructure provided by the Hyak supercomputer system at the University of Washington.
6 Supplemental Information
6.1 Physical Motivation for Features
6.1.1 DAMS
The DAMS are useful for discriminating between crystalline structures that include null or anomalously small QAMS. In Figure 11 the local structure in (a) is clearly anisotropic and should therefore produce a large SAMS signal. However, from (b) we can see that there is a situation for which the the QAMS can made to equal zero (for a given distortion of the ligands), specifically $\alpha=1/\sqrt{2}$.
This example highlights a flaw in the QAMS that is inherent to its construction, namely that it’s possible to produce a zero QAMS for a clearly anisotropic system. Here we use a relatively simple example, but the same result can be found for other geometries, and the condition will also depend on the normalization constant $m$. However, this failing can be offset by the inclusion of the DAMS, which can only be made to be zero for $\alpha=0$, and therefore provide orthogonal information to the model that can inform it local anisotropy that it would not be able to infer from the QAMS alone.
Figure 11: (a) Distorted square planar structure with a TM ion (silver) lying at (0, 0, 0) and four ligands lying along the $x$, $y$ axis that have all been translated positively by $\alpha$ along the $z$ axis. The charge $q_{n}$ of each ligand has been assigned as 1 for convienience. (b, c) The calculation of the (QAMS, DAMS) for an arbitrary normalization constant $m$.
6.2 Permutation Feature Importance
Permutation Feature Importance (PFI) begins with the definition of a performance metric. In our case we use the negative mean squared error (NMSE):
$\text{NMSE}(y,\hat{y})=-\frac{1}{n}\sum_{i=1}^{n}\big(y_{i}-\hat{y}_{i}\big)^{2},$
where $y_{i}$ are the true values, $\hat{y}_{i}$ are the predicted values, and $n$ is the number of samples. We adopt NMSE following the convention in scikit-learn, where higher scores indicate better performance. Since MSE is minimized, its negation ensures that values closer to zero represent better performance, while more negative values indicate worse performance [sklearn_permutation_importance].
Let $\text{NMSE}_{0}$ denote the baseline score of the model on the unpermuted dataset. For each feature $j$, we construct a permuted dataset $X^{\pi(j)}$ by randomly shuffling the entries in column $j$ while leaving all other columns unchanged. We then recompute the score
$\text{NMSE}_{j}=\text{NMSE}(y,\hat{y}^{\pi(j)}).$
The PFI for feature $j$ in a single permutation trial is defined as
$\text{PFI}_{j}=\text{NMSE}_{0}-\text{NMSE}_{j}.$
This process is repeated $R$ times with different random permutations, and the average importance is taken as
$\overline{\text{PFI}}_{j}=\frac{1}{R}\sum_{r=1}^{R}\Big(\text{NMSE}_{0}-\text{NMSE}_{j}^{(r)}\Big).$

### v2

@@ 5 Acknowledgments
JJK and JJR acknowledge support from the Theory Center for Materials and Energy Spectroscopies (TIMES) at SLAC funded by DOE BES Contract DE-AC02-76SF00515. CAC was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2140004. GTS was supported by funding from the U.S. Department of Energy in the Nuclear Energy University Program under Contract No. DE-NE0009158. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or the U.S. Department of Energy. This work was partly facilitated through the use of advanced computational infrastructure provided by the Hyak supercomputer system at the University of Washington.